OpenStudy (anonymous):

The average height of an apple tree is 4.00 meters. How long would it take an apple falling from that height to reach the ground? Given: g = -9.8 meters/second^2 -0.82 Seconds -1.0 Seconds -1.53 Seconds -2.3 Seconds

OpenStudy (theeric):

Hi! You need an equation for this, do you agree? It has to have distance and acceleration! Any ideas?

OpenStudy (anonymous):

Yes. I have none at all lol

OpenStudy (theeric):

\(d=v_it+\frac{1}{2}a\ t^2\) should suffice!

OpenStudy (theeric):

Wait...

OpenStudy (anonymous):

What? lol And what does the varables symbolize in the equation? Well, I know V, but what's D, A, and T?

OpenStudy (theeric):

Well, \(v_i=0\)

OpenStudy (theeric):

\(d\) is the distance, or the height \(a\) is the acceleration. Here, it's the acceleration due to gravity, which is usually \(9.8\), \(9.81\), or \(10\ m/s^2\) depending on your teacher. \(t\) is the time, how long it will take the apple to fall.

OpenStudy (theeric):

The equations is now \(d=\frac{1}{2}a\ t^2\).

OpenStudy (theeric):

The whole thing is\[x_f=x_i+v_i\ t+\frac{1}{2}a\ t^2\] The "\(_i\)" indicates "initial," and the "\(_f\)" indicates "final." They're initial and final properties across the interval. Here, the interval is the apple being on the tree to where the apple hits the ground.

OpenStudy (theeric):

And \(a\) for you is \(-9.8\ [m/s^2]\).

OpenStudy (theeric):

So you've found the answer? :) Or do you have any questions?

OpenStudy (theeric):

I don't see a choice for the answer I got.

OpenStudy (anonymous):

Actually, to be honest, I guessed and got it wrong :/ I'm only taking physics (trying to get at least a B) to get by and graduate with honors. But thanks anyways! For some reason, I'm just not understanding some criteria in Physics as greatly as I wish I could, especially since I'm self-teaching myself online -_- Again, thanks, theEric!

OpenStudy (anonymous):

Lol I get frustrated easily with things like this.

OpenStudy (theeric):

Oh gosh! Self teaching! Good luck! Did it give you the correct answer?

OpenStudy (theeric):

I got\[\sqrt{\frac{(2)(-4\ [m])}{-9.8\ [m/s^2]}}\]

OpenStudy (theeric):

\(\approx 0.903508\ [s]\)

OpenStudy (anonymous):

Thank you! :) I randomly clicked 0.82 and got it right.

OpenStudy (theeric):

Haha, nice! And I hope OpenStudy can help you out in the future :)

OpenStudy (theeric):

Take care!

OpenStudy (anonymous):

Thank you very much! You take care too :] Thanks for your time and effort.

OpenStudy (theeric):

My pleasure, thanks.